You have agreed to pay $200 three months from now and $500 six months from now in exchange for receiving $700 15 months fromhow to solve this using hp10bii calculator
Answer #1

Rate per 3 month or Quarterly rate (i)=8%/4=   0.02
Payment in 3 month (FV) =   -200
Number of Quartly period (n)= 3/3=   1
  
Payment in 6 month (FV)=   -500
Number of Quartly period (n)= 6/3=   2
  
Receipt in 15 month (FV)=   700
Number of Quartly period (n)= 15/3=   5
  
Net present value is sum of all PV. PV = FV/(1+I)^n  
(-200/(1+0.02)^1)+(-500/(1+0.02)^2)+(700/(1+0.02)^5)  
-42.65125511  
  
So Present Value of these cash flows is   -$42.65
  
Calculator function (for example 1 Payment)  
Rate (i/y)=   2%
Number of period (N)=   1
Payment per period (PMT)=   0
Future Value (FV)=   -200
  
Press CPT and PV, result is    -$196.08
  
Sum of all PV is PV of cash flows